### v2

# Differential pumping for kHz operation of a Laser Wakefield accelerator based on a continuously flowing Hydrogen gas jet

Abstract
Laser-Wakefield Accelerators (LWFA) running at kHz repetition rates hold great potential for applications. They typically operate with low-energy, highly compressed laser pulses focused in high-pressure gas targets. Experiments have shown that the best-quality electron beams are achieved using Hydrogen gas targets. However, continuous operation with Hydrogen requires a dedicated pumping system. This work presents the design of a differential pumping system, enabling, for the first time, continuous operation of our kHz LWFA using a high-pressure Hydrogen gas jet. The system successfully maintained a pressure below $3\times 10^{-4}\,$mbar, even with a free-flowing gas jet operating at 140 bar backing pressure. Numerical fluid dynamics and optical simulations were used to guide and validate the system’s design.
preprint: AIP/123-QED
I Introduction
Laser wakefield acceleration (LWFA) is a process that enables the generation and acceleration of electron beams to relativistic energies over very short distances Tajima and Dawson (1979). An ultrashort laser pulse is focused into a plasma, and drives a large density amplitude plasma wave in which plasma electrons can be trapped and accelerated. Since the plasma is already ionized, the electric field of the plasma wave is not subject to the breakdown limit that exists in conventional radiofrequency (RF) cavities. Thus the accelerating field can reach amplitudes up to 4 orders of magnitude higher than in conventional RF cavities: acceleration occurs over much shorter distances, paving the way for compact accelerators.
In LWFA, the laser first propagates through a vacuum before being focused into a plasma, where the acceleration process occurs. The plasma is generated as the laser ionizes the gas supplied by the target system. Various types of gas targets are used in LWFA. Each type is designed to balance the required plasma density while minimizing gas leakage, which increases the pressure in the vacuum chamber where the laser propagates. Here are the most commonly used target systems: (i) gas jets: a supersonic gas flow expands from a nozzle into the vacuum chamber Semushin and Malka (2001); Schmid and Veisz (2012), (ii) gas cells: the gas is enclosed in a cell, creating an interaction volume in which the laser propagates through two apertures at the cell’s entrance and exit Osterhoff et al. (2008). These apertures must be small enough to minimize gas leakage in the main chamber, but large enough to avoid ablation from the laser which would in turn negatively affect the laser performance. In order to mitigate the leaks, the gas cells operate at lower gas pressures, resulting in lower plasma densities compared to gas jets. Additionally, some gas cells utilize differential pumping, where the gas is removed from the cell simultaneously as it is introduced Kirchen et al. (2021); Drobniak et al. (2023). A last commonly used target system is the capillary discharge target, where the plasma is generated via a high-voltage discharge within a capillary filled with gas. There, the laser propagates through the guided mode of the plasma channel formed during plasma expansion and interaction with the capillary walls Leemans et al. (2006). This approach is relevant when long acceleration distances are needed (in the range of 111\,1cm), but the entrance capillary aperture can be damaged by high-intensity lasers. It is worth noticing that similar target systems are also being employed in the field of nuclear physics. Notably, some of their experiments make use of gas jets combined with differential pumping techniques, though these systems generally operate at much lower densities than those in the present study Ulbricht, Clausnitzer, and Graw (1972); Bittner, Kretschmer, and Schuster (1979); Treichel, Isenbügel, and Marquardt (1983); Shapira et al. (1985). Early-stage designs and implementations of such differential pumping systems have been explored in the frame of LWFA Bohlen et al. (2022).
Advances in laser technology, along with the development of the target systems, have led to tremendous progress in the laser-plasma accelerator performances. Today, LWFA driven by 100 TW to PW scale laser systems can produce electron beams with energies ranging from 100100100\,100MeV to GeV over a few centimeters Mangles et al. (2004); Geddes et al. (2004); Faure et al. (2004); Kim et al. (2017); Oubrerie et al. (2022a); Miao et al. (2022). To date, the highest energy ever reached is 8 GeV using plasma discharge technology Gonsalves et al. (2019).
While pushing the energy frontier is important for high-energy physics, another line of research focuses on the development of high repetition-rate (100 Hz - 1 kHz) LWFAs, which holds significant potential for enhancing stability and enabling the acquisition of large datasets, essential for robust statistical analysis. Until recently, kHz laser systems were restricted to a few millijoules per pulse. To fulfill the conditions for laser-wakefield acceleration using such energies, it is necessary to tightly focus and strongly compress the laser pulse, and to work with very high electron densities in the plasma on the order of $n_{e}\sim 10^{20}\,\textrm{cm}^{-3}$ Faure et al. (2018); Guénot et al. (2017); Gustas et al. (2018); Salehi et al. (2017). The high electron density imposes the use of free-flowing gas jets, which can provide such densities while refreshing the target quickly enough to handle kHz operation. The continuous operation of LWFA at 1 kHz has been demonstrated using Nitrogen Guénot et al. (2017); Rovige et al. (2020). However, recent experimental work shows that using light gases such as Hydrogen is beneficial to LWFA and that higher electron energy and better spatial beam properties can be reached Salehi et al. (2021).
Regarding vacuum conditions, utilizing a nitrogen plasma is convenient: each $\mathrm{N_{2}}$ molecule provides 10 electrons, enabling a backing pressure of only 20 to 30 bar in the gas jets to achieve the required plasma density. Additionally, heavy gases like nitrogen are easier to pump than light gases Gaede (1915); Becker (1966). However, ionization effects can significantly distort the laser pulse in $\mathrm{N_{2}}$ as different levels are ionizes at different positions within the laser pulse, causing a decrease in accelerator performance. To counter this, utilizing a light gas such as hydrogen is necessary Monzac et al. (2024). So far, kHz-scale experiments using an $\mathrm{H_{2}}$ plasma have been limited to burst mode operation as pumping $\mathrm{H_{2}}$ poses significant challenges Salehi et al. (2017, 2021). When a free-flowing gas jet of $\mathrm{H_{2}}$ is used at high pressure, it results in considerable gas loading in the vacuum chamber where the experiment is conducted. This phenomenon was observed by Salehi et al. Salehi et al. (2021), who reported a rapid pressure increase in the main chamber, rising from $2.6\times 10^{-2}\,$mbar to $2.0\times 10^{-1}\,$mbar in just 1 second when using a free-flowing gas jet with a $\sim 150\,\text{\textmu}$ µm FWHM aperture and a backing pressure of $\sim 30\,$bar. They operated their accelerator in burst mode and avoided gas loading in their experiments.
This paper describes the design, implementation, and testing of a differential pumping scheme for a continuously flowing gas nozzle used with light gases in a kHz laser-plasma accelerator. It is structured as follows: Section II introduces the physical principles underpinning the design of the differential pumping system, as well as a thorough analysis demonstrating its necessity and effectiveness. In Section III, we conduct fluid simulations and laser propagation simulations to verify the absence of laser distortion caused by gas stagnation in the small chamber. Section IV details the integration of this apparatus into our experimental setup and presents the results obtained through the implementation of the differential pumping system. Finally, Section V provides a summary of the findings and concludes the paper.
II Differential pumping with gas jets and high pressure light gases
The optimal plasma density and length for a laser-plasma accelerator depend on the physical properties of the laser pulse. In the case of a few mJ, few fs pulse, the plasma density should typically be $n_{e}\sim 10^{19}-10^{20}\,\mathrm{cm}^{-3}$ and the plasma length $L=100-200\,\text{\textmu}$ µm. To reach such plasma parameters, we use supersonic cylindrical gas nozzles and backing pressures ranging from a few tens of bar to about 150150150\,150bar. A cross-section diagram of the gas jet is shown in Figure 1b. In the following, we will refer to the gas nozzle parameters as follows: the diameter of the nozzle throat and nozzle opening are respectively written $D^{*}$ and $D$ and the corresponding throat and opening surfaces are $A^{*}$ and $A$, where we simply have $A=\pi D^{2}/4$. For example, a 60/180 gas nozzle has $D^{*}=60\,\text{\textmu}$ µm, $D=180\,\text{\textmu}$ µm, and provides the right plasma length and molecular density $n_{N_{2}}\sim 10^{19}\,\mathrm{cm}^{-3}$ when using $\mathrm{N_{2}}$ gas with pressure $P_{back}=30\,$bar backing pressure (see Figure 1a). The laser pulse is intense enough to ionize Nitrogen atoms to N5+, so that each $\mathrm{N_{2}}$ molecule releases 10 electrons, leading to the required plasma electron density of $n_{e}\sim 10^{20}\,\mathrm{cm}^{-3}$.
Figure 1: a) Electronic density $n_{e}$ of the plasma as a function of the backing pressure in the gas jet and the nature of the gas for a 60/180 supersonic nozzle. b) Cross section of the supersonic gas jet, $D^{*}$ and $D$ are respectively the diameter of the nozzle throat and nozzle opening.
II.1 Differential pumping use in high repetition rate LWFA
The pumping system is designed so that the residual equilibrium pressure in the experimental chamber, $P_{eq}^{(1)}$ stays sufficiently low and does not affect laser propagation. We set this limit empirically to $P_{eq}^{(1)}<10^{-2}\,$mbar since above this value, turbo-molecular pumps become inoperative, overheat, and the pumping system diverges. The residual pressure at equilibrium $P_{eq}^{(1)}$ in the chamber can be estimated knowing the pumping speed $R_{eff}[\mathrm{l/s}]$ and the gas leak rate due to the gas jet $Q[\mathrm{mbar\cdot l/s}]$ by simply writing that:
$P_{eq}^{(1)}[\mathrm{mbar}]=Q[\mathrm{mbar\cdot l/s}]\ /\ R_{eff}[\mathrm{l/s}]$ (1)
This allows to easily estimate the gas leak rate in our experiment. Typically, in Nitrogen, using a nozzle with $D^{*}=60\,\text{\textmu}$ µm and a backing pressure of $P_{back}=20\,$bar, we measure $P_{eq}^{(1)}=3\times 10^{-3}\,$mbar. Knowing that the pumping speed of our turbo-molecular pump is $R_{eff}=4000\,$l/s for Nitrogen, we obtain that the gas leak is $Q=12\,$mbar.l/s.
Let us now calculate what happens if one uses $\mathrm{H_{2}}$ instead of Nitrogen. We assume complete ionization: each $\mathrm{H_{2}}$ molecule now releases 2 electrons instead of 10 in $\mathrm{N_{2}}$. In addition, we use a 1D isentropic expansion model for compressible gas flows to extrapolate the molecular density in $\mathrm{H_{2}}$. The molecular density in the flow, $n_{mol}$, can be expressed according to the Boltzmann constant $k_{B}$, the Mach number $M_{Mach}$ and the initial pressure and temperature, $P_{0},\,T_{0}$ in the reservoir Zucker and Biblarz (2002):
$n_{mol}=\frac{P_{0}}{k_{B}T_{0}}\left(1+\frac{\gamma-1}{2}M_{Mach}^{2}\right)^{-\frac{1}{\gamma-1}}\$ (2)
The coefficient $\gamma=c_{p}/c_{v}$ is the ratio of specific heat at constant pressure $c_{p}$ over the specific heat at constant volume $c_{v}$. For perfect diatomic gases such as $\mathrm{N_{2}}$ and $\mathrm{H_{2}}$, $\gamma=7/5$, whereas for a perfect monoatomic gas, such as $\mathrm{He}$, $\gamma=5/3$. For our gas jet, where $M_{Mach}\gtrsim 3$, with $P_{0}=20\,$bar and $T_{0}=300\,$K, we obtain the same molecular density at the exit of the nozzle $n_{mol}=4.3\times 10^{18}\,\mathrm{cm}^{-3}$ for $\mathrm{N_{2}}$ and $\mathrm{H_{2}}$. The resulting electronic density is $n_{e}=4.3\times 10^{19}\,\mathrm{cm}^{-3}$ for $\mathrm{N_{2}}$ and $n_{e}=8.6\times 10^{18}\,\mathrm{cm}^{-3}$ for $\mathrm{H_{2}}$. The final electronic density as a function of the backing pressure for different gases is shown in Figure 1a. Using $\mathrm{H_{2}}$, the backing pressure should be increased by a factor of 5 to reach the same electronic density as in $\mathrm{N_{2}}$, i.e. $P_{0}=100\,$bar, which means a 5-fold increase in the gas flow. In addition, the leak rate depends on the mass of the molecule as $Q\propto 1/\sqrt{M}$ which is another aggravating factor for the case of $\mathrm{H_{2}}$ (see next section). Thus, for the same final electron plasma density, the leak rates in $\mathrm{H_{2}}$ and $\mathrm{N_{2}}$ are related by the Graham’s law of diffusion:
$Q_{H_{2}}=5Q_{N_{2}}\left(\frac{M_{N_{2}}}{M_{H_{2}}}\right)^{1/2}$ (3)
so that instead of $Q=12\,$mbar.l/s for 20 bar of $\mathrm{N_{2}}$, the leak rate would be $Q=224\,$mbar.l/s for 100 bar of $\mathrm{H_{2}}$. To compensate for that increase and keep the background pressure below $P_{eq}^{(1)}=3\times 10^{-3}\,$mbar, as when using $\mathrm{N_{2}}$, the pumping speed should be higher than $R_{eff}=7.4\times 10^{4}\,$l/s which is unrealistic as the largest turbo-molecular pumps on the market have maximum pumping speeds of 4000 l/s. Moreover, one should notice that $\mathrm{H_{2}}$ is more difficult to pump than $\mathrm{N_{2}}$. Overall, heavy gases are easier to pump than light gases; the compression ratio of the turbo-molecular pump goes as $K\propto\exp(\sqrt{M})$ with M the molar mass of the gas, and the pumping speed goes as $S\propto-\ln(K)$ Gaede (1915); Becker (1966), so the lighter the gas, the slower the pumping speed of the turbomolecular pump. This has been the bottleneck for using continuous flows of $\mathrm{He}$ or $\mathrm{H_{2}}$ in LWFA. For example, in our system, the pumping speed of our turbo-molecular pump drops down from 400040004000\,4000l/s in $\mathrm{N_{2}}$ to 300300300\,300l/s in $\mathrm{H_{2}}$ at $P=10^{-1}\,$mbar. In practice, this means that for $P_{0}=100\,$bar of $\mathrm{H_{2}}$, the equilibrium pressure would be around $P_{eq}^{(1)}\simeq 0.5\,$mbar, which is likely too high for the turbo-molecular pump to operate. Similar conclusions could be reached for $\mathrm{He}$.
Figure 2: Without differential pumping: equilibrium pressure in the main chamber versus backing pressure for a gas nozzle with throat diameter $D^{*}=100\,\text{\textmu}$ µm. The pumping system is a turbo-molecular pump 400040004000\,4000l/s.
This has been verified experimentally: we have measured the pressure in the vacuum chamber in which a supersonic nozzle lets out a continuous flow of gas. The vacuum chamber is pumped by a $R_{eff}=4000\,$l/s turbo-molecular pump connected to a primary pump with pumping speed 130130130\,130l/s. In Figure 2, we show the variation of the equilibrium pressure versus backing pressure for $\mathrm{N_{2}}$, and $\mathrm{He}$. The results show very clearly that for $\mathrm{He}$, the equilibrium pressure increases abruptly above $P_{back}=20\,$bar and the pumping system diverges above that, not being able to dispose of the high $\mathrm{He}$ flow. This clearly confirms that a specific solution should be implemented for pumping high-pressure ($P>100\,$bar) light gases.
II.2 Design of the differential pumping system
An alternative solution is to use a differential pumping system: the gas jet is enclosed into a small vacuum chamber, the differential chamber, that is pumped by a powerful primary pump able to work efficiently at the mbar level with a pumping speed of 360360360\,360l/s. The idea is to pump out most of the gas from the small chamber while tolerating a higher equilibrium pressure at the mbar level, $P_{eq}^{(2)}\sim\,$mbar within the differential chamber subsystem. Then, by restricting the leakage from the differential chamber to the main chamber, the pressure can be kept below $P_{eq}^{(1)}<10^{-3}\,$mbar in the main chamber. In this case, the laser propagation is not affected by the residual gas, provided that the mbar region in the differential chamber is sufficiently short, e.g. millimeter scale. Figure 3 shows a conceptual schematic of the differential scheme that is implemented in the experiment. The laser enters and exits the chamber via two cones that terminate with 111\,1mm holes for restricting gas leakage. In addition, the “high” density region is limited to a $4$-mm gap in between the two cones to limit its impact on laser propagation.
Figure 3: a) Schematic of the differential chamber. b) Conceptual design for differential pumping with a free-flowing high-density gas jet.
1D model used to design the differential pumping
Interestingly, the experimental measurement of the gas leak rate $Q$ can be compared to a theoretical estimation using well-known fluid gas dynamics Zucker and Biblarz (2002). An estimation of the mass flow rate can be obtained using a 1D isentropic model of the gas flow in the nozzle. For a 1D stationary and turbulent flow, the mass flow rate $\dot{m}=\rho\nu A$ is constant throughout the flow; at the nozzle throat, it reduces to:
$\dot{m}=\rho^{*}\nu^{*}A^{*}=\gamma^{1/2}\left(\frac{\gamma+1}{2}\right)^{-\frac{\gamma+1}{2(\gamma-1)}}\frac{P_{0}}{\sqrt{RT_{0}}}A^{*}$ (4)
where $\rho^{*},\,\nu^{*}$ and $A^{*}$ are respectively the mass density, fluid velocity, and the surface at the nozzle throat, $P_{0}$ and $T_{0}$ are the pressure and temperature in the reservoir. Note that here, subscript 0 refers to the reservoir, i.e. the stagnation volume where the fluid velocity is $\nu_{0}=0$. The leak rate in mbar.l/s can then be obtained from $Q\,=PdV/dt=k_{B}T\frac{dN}{dt}$ which can eventually be written as:
$Q=\dot{m}RT$ (5)
[…]
Figure 4: Equilibrium pressure in the chambers as a function of effective pumping speed of the differential chamber $R_{eff}$ for a $\mathrm{H_{2}}$ gas leak of 320320320\,320mbar.l/s, corresponding to a backing pressure of 150150150\,150bar and a nozzle with $D^{*}=60\,\text{\textmu}$ µm. The pumping speed for the main chamber is assumed to be 400040004000\,4000l/s and the two holes for the differential pumping are 111\,1mm in diameter.
Using this model, it is straightforward to determine $P_{eq}^{(1)}$ and $P_{eq}^{(2)}$, respectively the equilibrium pressures in the experimental and differential chambers, that are reached as a function of the pumping speed $R_{eff}$ used for pumping the differential chamber. Figure 4 shows the results for a backing pressure of $P=150\,$bar in a nozzle with throat diameter $D^{*}=60\,\text{\textmu}$ µm. It is clear that the effective pumping speed needs to be larger than $R_{eff}>100\,$l/s to maintain the pressure in the differential chamber at the mbar level. When this condition is fulfilled, these estimations confirm that the equilibrium pressure in the main chamber remains at the level $P_{eq}^{(1)}\sim 10^{-5}\,$mbar, when the turbo-molecular pump is running at 4000 l/s. Here, we use the concept of effective pumping speed $R_{eff}$, where $R_{eff}$ is usually smaller than the nominal pumping speed $R_{nom}$ due to the finite conductance of the various elements of the pumping system (tubes, bellows, etc…) and also because the pumping speed varies with the operating pressure. This has to be carefully taken into account when designing the pumping system.
Once the differential pumping system had been set up, we measured the variation of the equilibrium pressure versus backing pressure for $\mathrm{N_{2}}$ and $\mathrm{H_{2}}$. Figure 5 shows the measurements: with the differential pumping apparatus, the pressure in the main chamber $P_{eq}^{(1)}$ remains below $3\times 10^{-4}\,$mbar, even with a backing pressure of 140140140\,140bar of $\mathrm{H_{2}}$ in the gas jet.
Figure 5: With differential pumping: equilibrium pressure in the main chamber versus backing pressure for a gas nozzle with throat diameter $D^{*}=60\,\text{\textmu}$ µm. The differential chamber is now pumped by a primary pump of 360360360\,360l/s.
III Numerical simulations
III.1 Fluid dynamics simulations
In order to obtain more accurate estimates of the gas flow in the nozzle and the differential chamber, we have performed numerical simulations using the computational fluid dynamics software ANSYS Fluent. Simulations were performed using molecular hydrogen modeled by the Aungier-Redlich-Kwong real gas equation. 3D geometry was used. The results are shown in Figure 6 for $\mathrm{H_{2}}$ with a backing pressure $P_{back}=100\,$bar and a nozzle throat of $D^{*}=60\,\text{\textmu}$ µm. In the simulations, the boundary conditions at the entrance of the cones and the pumping outlet were taken to be $P=0$. The simulation results shown in Figure 6a confirm that the pressure remains at the mbar level in the differential chamber, away from the region where the high-density jet is present. Figure 6b shows a dramatic decrease of the density by three orders of magnitude in the gap region between the two cones. The density then steadily decreases along the cone which is favorable for coupling the intense laser into the high-density gas jet without prior interaction. Thus, the gas remains well-confined within the differential chamber, but there is still some stagnation observed about 1 cm before the gas jet, potentially impacting laser performance.
Figure 6: a) Results of CFD simulations showing the pressure in the volume of the differential chamber. b) Lineout of the molecular density along the laser axis.
III.2 Propagation of the laser pulse up to the gas jet
Figure 7: Evolution of the laser parameters as it propagates through the gas profile from $z=-10\,$ mm to $z=-200\,\text{\textmu}$ µm. a) Evolution of the laser intensity $I$ in $\mathrm{N}$, $\mathrm{He}$ and $\mathrm{H}$. b) Evolution of the waist of the pulse $w_{0}$ in the three gases. c) Electron density along the propagation in the three gases and the target electron density (corresponding to fully ionized $\mathrm{He}$ or $\mathrm{H}$ gas, and up to the $5^{\mathrm{th}}$ in $\mathrm{N}$ gas).
To evaluate whether the stagnating gas affects laser propagation, we have performed simulations of the propagation of the laser pulse from the parabola (at $z=-10\,$cm) to the jet (at $z=-200\,\text{\textmu}$ µm) using an optical propagation code. The simulations were conducted using three gases : $\mathrm{H}$, $\mathrm{He}$ and $\mathrm{N}$ in order to assess the potential impact of different gases on the laser pulse. The propagation of the laser through an initially neutral gas was modeled using the open-source library “Axiprop” Andriyash (2024). This library uses various optical propagators, depending on the specific conditions (gas or vacuum for instance) and concatenate them to handle more complex scenarios. For our study, the paraxial propagator was used for vacuum propagation from the parabola ($z=-10\,$cm) to the point where the gas density becomes non-negligible ($z=-10\,$mm), as it handles drastic changes in beam size Goodman (2005). For the region with gas ($z=-10\,$mm to $z=-200\,\text{\textmu}$ µm), a non-paraxial propagator was applied, following the method by Oubrerie and al. Oubrerie et al. (2022b). Both propagators were implemented using Hankel transforms to account for cylindrical symmetry Oubrerie et al. (2022b). To model the propagation of the laser in the gas/plasma, the solver that propagates the field within the plasma is similar to the one described by Couairon and al. Couairon et al. (2011). Ionization of the gas is simulated using the ADK model Ammosov, Delone, and Krainov (1986); Augst et al. (1989). In the gas region, grid sizes were set to $\Delta z=5.0\,\text{\textmu}$ µm for $\mathrm{He}$ and $\mathrm{N}$, $\Delta z=4\,\text{\textmu}$ µm for $\mathrm{H}$, and $\Delta r=0.64\,\text{\textmu}$ µm in all three gases.
[…]
Figure 7 illustrates the evolution of the laser intensity ($I$) and waist ($w_{0}$) along its propagation through the density profiles of $\mathrm{N}$, $\mathrm{He}$, and $\mathrm{H}$ gases, with vacuum included for reference. Overall, the intensity and beam size remain nearly identical across vacuum and gas media, indicating that the effect of residual gas is largely negligible. However, at the entrance of the gas jet, where gas density inevitably rises, small differences become noticeable, causing slight perturbations in the laser pulse. Figure 7a shows a slight reduction in laser intensity when propagating through gas compared to vacuum. This reduction is most pronounced in $\mathrm{H}$ (reduction about 17%), and there is a smaller drop of roughly 8% in $\mathrm{N}$ and $\mathrm{He}$. Figure 7b highlights a slight increase in beam waist due to ionization-induced defocusing, with an increase of about 7% in $\mathrm{H}$, 5% in $\mathrm{N}$, and 4% in $\mathrm{He}$ relative to vacuum. Finally, Figure 7c shows that $\mathrm{H}$ fully ionizes early in the propagation, around $z=-4.4\,$mm, while $\mathrm{He}$ and $\mathrm{N}$ reach full ionization closer to $z=-700\,\text{\textmu}$ µm. These slight differences explain the laser pulse’s more noticeable degradation after propagating through $\mathrm{H}$, compared to $\mathrm{N}$ or $\mathrm{He}$.
Figure 8: Spectral evolution of the laser pulse after propagation in vacuum, and initially neutral $\mathrm{N}$, $\mathrm{He}$ and $\mathrm{H}$ gases.
Also, the interaction of the pulse with all three gases shifts the spectrum towards the blue. Figure 8 shows the spectrum of the initial pulse and of the pulse at the end of the simulations, integrated over the radial axis, after propagating through vacuum, $\mathrm{N}$, $\mathrm{He}$, or $\mathrm{H}$. When propagating through the three gases, the spectrum shifts slightly to higher frequencies, with a maximum normalized frequency shift of $\omega_{0,\text{max}}/\omega_{0}=1.02$. The laser pulse behaves similarly in all three gases, with no significant differences regarding the final spectrum.
The stagnating gas has a minimal impact on laser propagation. While the interaction of the laser pulse with the gases results in a slight reduction in intensity, some increase in beam size, and a minor blueshift in the spectrum, these changes are negligible. The effects are slightly more pronounced in $\mathrm{H}$ due to its lower ionization threshold and the fact that it fully ionizes significantly earlier than $\mathrm{He}$ and $\mathrm{N}$.
Figure 9: Experimental results obtained in a $\mathrm{H_{2}}$ (doped $\mathrm{N_{2}}$) plasma. (a) Single-shot beam profile samples obtained over 10 consecutive shots with $n_{e}=1.1\times 10^{20}\,\mathrm{cm}^{-3}$. (b) Electron spectra. The thickness of the lines corresponds to the RMS fluctuations of the spectrum. (c) Stability over a 100 shots with $n_{e}=7.4\times 10^{19}\,\mathrm{cm}^{-3}$. (i) Charge, (ii) spectra and (iii) beam profile in the $x$ direction, for each shot a slice of 3 pixels around the centroid is shown.
IV Implementation of the differential pumping apparatus : experimental results
After designing and simulating the differential pumping system, we implemented it in our experimental setup to test it in situ, aiming to explore the performance potential of LWFA at 1kHz with this system.
IV.1 Experimental set-up
[…]
Specifically, we used shocked gas jets and Hydrogen doped with 2% of Nitrogen. Shocked jets helped confine the electron-trapping region, resulting in electron beams with reduced energy spreads and improved stability Bulanov et al. (1998); Suk et al. (2001); Tomassini et al. (2003); Kim, Hafz, and Suk (2004). Additionally, shock-assisted ionization injection demonstrated an increase in trapped charge compared to pure shock injection while maintaining excellent stability properties Thaury et al. (2015).
The electron beam charge, spatial profile, and beam pointing are measured using a calibrated YAG screen imaged onto a 14 bits CCD camera. The electron energy spectrum is measured onto the same YAG screen, by inserting a motorized magnetic spectrometer in the beam, which relies on a set of permanent magnetic dipoles creating a magnetic field of 0.120.120.12\,0.12T on a 202020\,20mm length. Each image presented here was obtained by acquiring data over 111\,1ms (so each image corresponds to a single shot). For each plasma density studied, a series of 100 acquisitions is taken. Statistics over electron beam parameters are then obtained by averaging over the acquisition series, and the uncertainties represent the RMS deviation from the mean value.
IV.2 Results
Table 1: Stability of the beam : RMS fluctuations of the key beam parameters.
$n_{e}\ (\mathrm{cm}^{-3})$ $7.4\times 10^{19}$ $1.1\times 10^{20}$
Point. stab. x 1.46 mrad rms 1.04 mrad rms
Point. stab. y 1.97 mrad rms 2.26 mrad rms
$Q$ 1.2 pC 3.5 pC
(±$\pm 6.0\,\%$ rms) (±$\pm 6.7\,\%$ rms)
$E_{mean}$ 5.02 MeV 3.3 MeV
(±$\pm 2.4\,\%$ rms) (±$\pm 1.8\,\%$ rms)
$E_{spread}$ 0.99 MeV 1.44 MeV
(±$\pm 26\,\%$ rms) (±$\pm 6.9\,\%$ rms)
$\theta_{x}$ 13.0 mrad 18.2 mrad
(±$\pm 10\,\%$ rms) (±$\pm 13\,\%$ rms)
$\theta_{y}$ 12.1 mrad 16.2 mrad
(±$\pm 14\,\%$ rms) (±$\pm 11\,\%$ rms)
For each measurement, we optimized the laser pulse duration and focal position. We explored densities ranging from $n_{e}=7.4\times 10^{19}\,\mathrm{cm}^{-3}$ to $n_{e}=1.2\times 10^{20}\,\mathrm{cm}^{-3}$. Figure 9a shows a typical electron beam profile at the higher density of $n_{e}=1.2\times 10^{20}\,\mathrm{cm}^{-3}$. The beam exhibits a quasi-circular shape with minimal divergence, with very good stability shot to shot as shown in Table 1. Figure 9b shows the corresponding electron spectra for two different densities, averaged over 100 shots. The thickness of the lines corresponds to the RMS fluctuations in the dataset. Both spectra reveal low fluctuations, with well-defined energy peaks, confirming the consistent beam quality. The stability of the beam is further demonstrated in Figure 9c, where we present 100 consecutive single-shot acquisitions. The low variance in beam parameters such as charge, divergence, and energy indicates the high level of stability achieved with this setup. Table 1 summarizes the relative fluctuations of key beam parameters for the two plasma densities. Notably, the beam-pointing fluctuations remain below 2.3 mrad RMS, with charge fluctuations less than 7%, and energy fluctuations as low as 2% shot-to-shot, confirming excellent beam stability throughout the experiment.
V Conclusion
In this paper, we introduced a novel apparatus that combines high-pressure gas jets with a differential pumping system, enabling the continuous operation of a kHz laser-wakefield accelerator utilizing hydrogen plasma generated from a free-flowing gas jet.
The theoretical calculations demonstrate the importance of this system in meeting operational requirements and guide the design process. The chamber pressure was successfully maintained below $3\times 10^{-4}\,$mbar, even with a backing pressure of up to 140 bar of hydrogen in the 180 µm aperture gas jet, confirming the effectiveness of this vacuum system. Then, fluid simulations demonstrated that the gas behavior within the chamber does not adversely affect laser pulse propagation.
Finally, we conducted laser wakefield acceleration experiments to test the efficacy of the differential pumping setup, yielding highly stable and well-defined electron beams. This accomplishment represents the first continuous operation at kHz repetition rates with hydrogen, thereby marking a significant advancement in kHz LWFA using light gases.
Acknowledgements.
This project has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement JRA PRISE no. 871124 Laserlab-Europe and IFAST under Grant Agreement No 101004730. This project was also funded by the Agence Nationale de la Recherche under Contract No. ANR-20-CE92-0043-01, and benefited from the support of Institut Pierre Lamoure via the chair “Accélération laser-plasma haute cadence”.
This work has benefited from a grant managed by the Agence Nationale de la Recherche (ANR), as part of the program ‘Investissements d’Avenir’ under the reference (ANR-18-EURE-0014).
Financial support from the Région Ile-de-France (under Contract No. SESAME-2012-ATTOLITE) and the Extreme Light Infrastructure-Hungary Non-Profit Ltd (under Contract No. NLO3.6LOA) is gratefully acknowledged. We also acknowledge Laserlab-Europe, Grant No. H2020 EC-GA 654148 and the Lithuanian Research Council under Grant agreement No. S-MIP-21-3.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.